OpenStudy (abbycross167):

Can someone please check my work? Please? I would really appreciate it... Question~ Solve the systems using the substitution method. y = (2/3)x + 1 y + x = 6

OpenStudy (abbycross167):

My Answer~ Solve one equation for either variable ~ Y+x=6 In the other equation, replace the variable from S1 with what it equaled (2/3)x+1+x=6 Solve the equation in step 2 for the other variable: Multiply both sides by 3 to clear the fraction~ 3(2/3)x+1+x=6*3 Combine like terms~ 2x+3+3x=18 Subtract 3 from both sides~ 5x+3-3=18-3 Solve for x ~ 5x=15 Write your answer as an ordered pair (x, y). (3,3)

OpenStudy (abbycross167):

What my teacher said was wrong~ What is this? Solve one equation for either variable ~ Y+x=6 = the rest of your steps don't make sense to me. can you help me?

Mehek (mehek14):

you got the answer correct. . .

OpenStudy (phi):

You start by saying Solve one equation for either variable then you write down the 2nd equation. But I would write down the first equation y = (2/3)x + 1 because "solve for a variable" means have the variable by itself on one side of the = sign. In the first equation y is by itself on the left side. In other words, the first equation is already "solved for y". So I would start by saying 1. Solve one equation for either variable. We use equation 1 because it is already solved for y: y = (2/3)x + 1

OpenStudy (phi):

Then I would say 2. In the other equation, y+x= 6 replace the variable y with what it equaled. The new equation is (2/3)x+1+x=6

OpenStudy (phi):

Next step, the easiest thing to do is add -1 to both sides 3. simplify the equation by adding -1 to both sides. It is (⅔) x + x = 5 next, you can multiply both sides by 3 (to get rid of the fraction) or you can add ⅔ x + 3/3 x = 5/3 x But let's multiply by 3: 4. Multiply both sides by 3 2x+3x = 15 5. combine like terms 5x= 15 6 divide both sides by 5 x=15/5 x= 3 7. Use either equation to solve for y. the first equation, with x=3 is y = (⅔)*3 + 1 y= 2+1 y=3

OpenStudy (phi):

Your write-up is mostly ok. Did your teacher not like it ?

OpenStudy (abbycross167):

I guess she didn't, but I have done many other answers like the one I posted... Thank you so much for showing me how I should write the steps!! You're amazing!!